OpenStudy (anonymous):

Betty is thinking of 2 consecutive integers whose sum is 41. Let x represent the smaller unknown integer. Write an equation showing that the sum of the two unknown integers is 41.

OpenStudy (anonymous):

so if two numbers are consecutive, you can write the first one as x and the next one as x+1. so X+X+1=41 and you can simplify that to 2X+1=41. then just solve for x by subtracting 1 and then dividing by 2. And you get x=20 and X+1=21.